find the absolute maximum and absolute minimum values of f in the given interval. Give exact values using radicals, as necessary.

f(t)=t-\sqrt[3]{t}, [-1,7]
Answer #1

i. f(t) = 0 (ors of Calculate any critical points s (t) = a (x - 35t) sinf 2013 s(t) = 1 = 34 2/3 213 3 2+213 - 343 t 1 =0